OpenStudy (zzr0ck3r):

\(g:[0,1]\rightarrow \mathbb{R}\) is continuous, \(g(0)>0\) and \(g(1)=0\). Show that there is an \(a\in (0,1]\) s.t. \(g(a) = 0\) and \(g(x)>0\) for all \(0\le x < a\) .

OpenStudy (zzr0ck3r):

A student brought this to the tutoring center where I work, and the problem is in a section that comes right before the intermediate value theorem, it is in the section for extreme value theorem. So they basically have the definition of continuity and the E.V.T . I would love some help to find a not to intense solution. It is from a advanced calculus class.

OpenStudy (anonymous):

Well, IVT would say what exactly? That \(b\in (0,1]\) where \(0 \leq g(b)<0\)?

OpenStudy (zzr0ck3r):

that there exists a x in (0,1) s.t. f(x) = 0.

OpenStudy (zzr0ck3r):

my real point about the IVT is that we cant use it.

OpenStudy (anonymous):

I thought IVT was something like \[ c \in (a,b) \implies f(c) \in(f(a), f(b)) \]or \((f(b), f(a))\) depending on which interval is smaller.

OpenStudy (anonymous):

not which interval is smaller, I mean depending on how you'd order the endpoints.

OpenStudy (zzr0ck3r):

This tablet is killing me, brb

OpenStudy (zzr0ck3r):

if \(f:[a,b]\rightarrow \mathbb{R}\) is continuous. Let \(x\) be in \((f(a),f(b))\), then there exists \(x_0\) in \((a,b)\) s.t. \(f(x_0)=c\)

OpenStudy (zzr0ck3r):

that c should be an x

OpenStudy (zzr0ck3r):

But anyway, I cant use it...

OpenStudy (anonymous):

why is that?

OpenStudy (zzr0ck3r):

Because it is in a section in the book that comes before IVT. So they don't have that tool.

OpenStudy (anonymous):

Oh, they just have EVT?

OpenStudy (zzr0ck3r):

I can use EVT and continuity, and anything about sequences.

OpenStudy (zzr0ck3r):

Correct. I was thinking if I could get the function down to an interval where we only have one zero, then we could take the abs value of the function and find the min. (assuming f(x) is not always positive for x not equal to 1).

OpenStudy (anonymous):

We know there is a minimum on \(g\), that minimum can't be at \(0\) because \(g(0) > g(1)\).

OpenStudy (anonymous):

I mean \(g(0)\) is not the minimum.

OpenStudy (anonymous):

Now it is possible that \(g(1)\) is the minimum, which means \(a=1\).

OpenStudy (anonymous):

If it isn't, then we would say there is some \(a\neq 1\) where \(g(a)\leq 0\)

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

or actually let's use \(b\) as not to confuse with the \(a\) we claim to exist.

OpenStudy (anonymous):

I think in the case where \(g(1)\) is not the minimum, then we will change the interval to \([0,b]\).

OpenStudy (zzr0ck3r):

lets just assume that its not always positive outside of g(1)

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

This is a bit complicated without intermediate value theorem isn't it.

OpenStudy (zzr0ck3r):

Yeah, apparently the student brought it to his prof and the prof could not solve it.

OpenStudy (anonymous):

I'm thinking we would take the limit like... \[ \lim_{x\to b^+}g(x) \]

OpenStudy (anonymous):

We know this limit exists and equals \(g(b)\) because it is continuous.

OpenStudy (anonymous):

In terms of epsilon delta, what does that tell us?

OpenStudy (anonymous):

Is it something like:\[ x-b<\delta \implies |g(x) - g(b)|<\epsilon \]

OpenStudy (anonymous):

Hmm, I think I should have put \(x-b>\delta\).

OpenStudy (anonymous):

I want it to be a limit going toward the positive direction.

OpenStudy (anonymous):

Uhg. I actually have two issues with this problem.

OpenStudy (anonymous):

The first is that it could hit 0 and infinite amount of times.

OpenStudy (zzr0ck3r):

yeah, I agree. So its like there is some 'first" time on a conceivably uncountable list.

OpenStudy (anonymous):

The second is trying to show that if the minimum is negative, we would want to throw it out while maintaining a closed interval.

OpenStudy (anonymous):

Like consider:

OpenStudy (anonymous):

I think we can work with this case a bit...